OpenStudy (anonymous):

f(x) = x3 + 3x2 - x - 3 Part A: What are the factors of f(x)? Show your work. How do I factor this when the 3 doesn't have any factors and the x has no coefficient?

OpenStudy (anonymous):

f(x) = x2(x+3) - 1(x+3)

OpenStudy (anonymous):

f(x) = (x2 -1) (x+3)

OpenStudy (anonymous):

to do these first you want to regroup the 1st and 2nd term then regroup the 3rd and fourth term

OpenStudy (anonymous):

I knew that, forgot about the invisible one next to x thnks :)

OpenStudy (anonymous):

then factor out the x^2 and -1 then you can rewrite it as (x^2 -1)(x+3)

OpenStudy (anonymous):

you're welcomed :)

OpenStudy (anonymous):

Could you help me out with the second and third part?

OpenStudy (anonymous):

sure, the second step that I do this problem?

OpenStudy (anonymous):

or another question?

OpenStudy (anonymous):

Part B: What are the zeros of f(x)? Show your work. but the zeros are just -1 and 3 for this?

OpenStudy (anonymous):

set x2-1 = 0 now find x set x+3 = 0 find x so to do these, set factors of the equation equate to 0.

OpenStudy (anonymous):

nope, x2 = 1 when you square root 1 , you get -1 AND 1

OpenStudy (anonymous):

1 squared = 1 -1 squared = 1

OpenStudy (anonymous):

Oh so 1 and -3 x^2 -1 =0 ---> x^2 = 1 and same with second

OpenStudy (anonymous):

so (1,0) and (-3, 0)

OpenStudy (anonymous):

(-1,0) (1,0) (3,0)

OpenStudy (anonymous):

*** (-3,0) I mean not (3,0)

OpenStudy (anonymous):

\[x^{2} = 1\] x= 1, -1 because \[(-1) ^{2} = 1 \] and \[1^{2} = 1\]

OpenStudy (anonymous):

Could you explain how you got 3? I'm still confused on that part.

OpenStudy (anonymous):

I said (-3,0) not (3,0) :) it was a typo from before CONCLUSION (-1,0)(1,0) (-3,0)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

art C: What are the steps you would follow to graph f(x)? Describe the end behavior of the graph of f(x). I think i got this down except I don't know if graph going left or right

OpenStudy (anonymous):

use the 0s to graph it, then end behavior starts from down to up

OpenStudy (anonymous):



OpenStudy (anonymous):

just plot the 0s first

OpenStudy (anonymous):

Ok I think I'm good now thanks for all help.

OpenStudy (anonymous):

glad it helps :)